QUESTION 15 6.25 points Save Answe An analyst has developed the following forecast for the overall economy and the returns for a particular stock State Boom Moderate Growth Flat Probability .25 .40 .15 Return .35 .15 .08
Bust 20 -.25 Find the coefficient of variation O 2.6642 1.8513 O.5402 O 2027
Answer #1

OPTION B

Mean=0.25*0.35+0.40*0.15+0.15*0.08+0.20*(-0.25)=0.1095

Coefficient of Variation=sqrt(0.25*(0.35-0.1095)^2+0.40*(0.15-0.1095)^2+0.15*(0.08-0.1095)^2+0.20*(-0.25-0.1095)^2)/(0.1095)=1.851309596